### v1

# Signal Detection in Near-field Communication with Unknown Noise Characteristics: A Diffusion Model Method

Abstract
In this letter, we present a diffusion model method for signal detection in near-field communication with unknown noise characteristics. We consider an uplink transmission of a near-filed MIMO communication system consisting of multiple mobile terminals and one base station with multiple antennas. Then, we proposed a Maximum Likelihood Estimation Diffusion Detector (MLEDD) aiming at learning the distribution of unknown noise. To this end, we define an error function via Bayes’ theorem to detect the source signal. Moreover, we present an implementation of the proposed framework. The performance of the proposed method in terms of bit error rate shows that it outperforms the MLE detector, Detection Network (DetNet), and Maximum Normalizing Flow Estimate method (MANFE) across different signal-to-noise ratios and noise distributions. Especially when the noise distribution is intractable, diffusion, as a state-of-the-art probability model, has the best distribution learning ability compared to other models. These results affirm that this framework can effectively detect signals in near-field scenarios.
Index Terms:
Diffusion model, signal detection, near-field communication, unknown noise statistics, generative models.
I Introduction
The transition from 5G to 6G wireless networks signifies a shift to a more sophisticated, immersive, and integrated network ecosystem with unprecedented speeds, low latency, and support for advanced technologies including artificial intelligence [1]. Among these technologies, large antenna arrays and extremely high-frequency bands significantly increase the Rayleigh distance [2], leading to non-negligible near-field electromagnetic (EM) effects. In contrast to far-field communication, which predominantly utilizes planar wave models, near-field communication necessitates precisely modeling spherical waves [1]. In near-field regions, the interaction among multiple devices creates more complex interference patterns that significantly affect data transmission and more intricate noise effects. Due to the complexities inherent, an optimized signal detection technique is essential for accurately estimating and extracting transmitted data , particularly when noise characteristics are unknown [3].
Signal detection is a classical problem that can be solved using model-based or learning-based algorithms. The linear detector, including the matched filter and the minimum mean squared error (MMSE) detector, is the most commonly used model-based detector [4]. When assuming the distribution of noise is tractable, such as additive white Gaussian noise (AWGN), the near-optimal detected signal can be obtained by linear detectors. Several advanced model-based detectors, such as the generalized approximate message passing method [5], can handle arbitrary noise. However, they still require approximate methods to compute the marginal posterior. On the other hand, Detection Network (Detnet) is a typical learning-based approach, which leverages the supervised learning framework to predict source signals [6]. Due to the deterministic nature of supervised learning, DetNet struggles to make accurate predictions when faced with the received signal mixed with complex noise. Generative models have recently gained popularity as probabilistic models in unsupervised learning. Via learning distribution rather than specific labels, they have numerous applications in physical layer communications, particularly in complex noise scenarios [7]. For instance, the Maximum Normalizing Flow Estimate method (MANFE) utilizes normalizing flow to detect MIMO signals, demonstrating excellent performance in scenes with unknown noise [8]. However, it cannot guarantee performance in complex scenarios such as near-field communication. Recently, the denoising diffusion model, inspired by non-equilibrium thermodynamics, has demonstrated its powerful ability to learn from unknown distributions in various tasks better than other generative models [9]. This superiority makes it a promising method for estimating complex noise distributions.
This letter proposes a Maximum Likelihood Estimation Diffusion Detector (MLEDD) to explore near-field signal detection with unknown noise characteristics. We consider an uplink transmission of a near-filed MIMO communication system and utilize the diffusion model to learn the unknown noise distribution. Then, we define an error function as an evaluation indicator to detect source signals. To the best of our knowledge, this is the first attempt to use generative models for signal detection in near-field communication scenarios. Numerical results demonstrate that the MLEDD with the same magnitude of complexity can achieve a better bit error rate (BER) performance than that of the traditional MLE detector, DetNet, and MANFE across various signal-to-noise ratios (SNR) and noise distributions.
II Preliminaries
II-A Near-field MIMO System Model
In this letter, we consider an uplink transmission of a near-filed MIMO communication system consisting of $M$ mobile terminals (MTs) and a base station (BS) equipped with $N$-antenna uniform linear array (Fig. 1). Denoting $r_{m}$ as the distance between the $m$-th MT and the center of the receive array antenna, the MT is located in the Fresnel zone (radiative near field) of the array, i.e., $0.62(D^{3}/\lambda)^{1/2}<r_{m}<2D^{2}/\lambda$, where $m=1,2,\ldots,M$, $D$ denotes array aperture and $\lambda$ is the carrier’s wavelength [2]. Assuming each MT transmits the source signal to the BS through independent paths, the received signal $\bm{y}$ can be written as follows:
$\bm{y}=\bm{A}\bm{x}+\bm{n},$ (1)
where $\bm{y}=[y_{1},\ldots,y_{N}]^{T}\in\mathbb{C}^{N\times 1}$, $\bm{x}=[x_{1},\ldots,x_{M}]^{T}\in\mathbb{C}^{M\times 1}$ represents the source signal vector, $\bm{A}\in\mathbb{C}^{N\times M}$ denotes the channel matrix, and $\bm{n}\in\mathbb{C}^{N\times 1}$ is an additive measurement noise, which is independent and identically distributed (i.i.d) with an unknown distribution $p_{n}(\cdot)$. We consider line-of-sight propagation in the small-coverage system and do not consider the multipath effect. In near-field communication, the channel matrix $A$ is composed of a series of array steering vectors, where each steering vector is derived from the accurate spherical wave. Specifically, the array steering vector of the $m$-the MT can be expressed as follows [10]:
$\begin{split}\bm{a}_{m}(\theta_{m},r_{m})=&[\kappa_{1,m}\exp\{-jk_{c}(r_{1,m}-r_{m})\},\ldots,\\ &~{}~{}~{}~{}~{}~{}~{}~{}\kappa_{N,m}\exp\{-jk_{c}(r_{N,m}-r_{m})\}],\end{split}$ (2)
where $\theta_{m}\in(-1,1)$ denotes the corresponding spatial angle from the $m$-th MT source to the receiver; $k_{c}=\frac{2\pi}{\lambda}$ represents the wave number; $\kappa_{n,m}=\frac{r_{m}}{r_{n,m}}$ is the corresponding amplitude; $r_{n,m}$ denotes the distance from the $n$-the array element to the $m$-the MT. According to the geometric relationship, we have
$r_{n,m}=\sqrt{r_{m}^{2}+\delta_{n}^{2}d^{2}-2r_{m}\delta_{n}\theta_{m}d},$ (3)
where $\delta_{n}=\frac{2n-N+1}{2},n=0,1,\ldots,N-1$, $d$ is the element spacing of the antenna array, and the coordinate of the $n$-the antenna is denoted as $(0,\delta_{n}d)$.
Figure 1: Illustration of the near-field region and the far-field region divided by the Rayleigh distance, with spherical wave and planar wave, respectively.
II-B Signal Detection
Generally, finding a formal solution via the MMSE in the near-field communication system with unknown noise characteristics is challenging since we do not have any prior knowledge of the noise. Thus, we adopt the maximum a posteriori (MAP) estimation to compute the detected solution, which can be written as follows:
$\begin{split}\widehat{\bm{x}}_{MAP}=\arg\max_{\bm{x}\in\mathcal{X}}p(\bm{x}|\bm{y})=\arg\max_{\bm{x}\in\mathcal{X}}p(\bm{y}|\bm{x})p(\bm{x}),\end{split}$ (4)
where $\mathcal{X}$ represents the set of all possible signal vectors, and $p(\cdot)$ denotes probability density function. Particularly, since there is no prior knowledge of the transmitted signal $\bm{x}$, the MAP estimation is equivalent to the maximum likelihood estimation (MLE), which can be formulated follows [8]:
$\begin{split}\widehat{\bm{x}}_{MAP}=\widehat{\bm{x}}_{MLE}&=\arg\max_{\bm{x}\in\mathcal{X}}p(\bm{y}|\bm{x})\\ &=\arg\max_{\bm{x}\in\mathcal{X}}p_{n}(\bm{y}-\bm{Ax}).\end{split}$ (5)
Based on Eq. (5), the signal detection problem under unknown noise characteristics requires us to accurately estimate the distribution of unknown noise and its MLE. However, model-based numerical methods require prior knowledge of the noise distribution, while traditional supervised learning AI models struggle with uncertain data, which leads to an increase in signal detection error [8].
II-C Denoising Diffusion Model
To accurately estimate the unknown noise distribution, we intend to use the diffusion model (DM), a generative model that excels in many complex distribution generation tasks, to detect the signal [9]. First, we briefly review the fundamental concepts, formulations, and training framework of DMs.
The DMs consist of two main processes: the forward diffusion process and the denoising process [11]. The forward diffusion process adds noise to the data, while the denoising process removes the noise to restore the data. Specifically, consider a data distribution $\bm{n}_{0}\sim p(\bm{n})$, the forward diffusion process adds $T$ Gaussian noises hierarchically to make the original distribution $p(\bm{n})$ converge to the standard normal distribution. Formally, The disturbance data $\bm{x}_{t}$ with $t$ times Gaussian noise added is expressed as,
$\begin{split}\bm{n}_{t}&=\sqrt{\alpha_{t}}\bm{n}_{t-1}+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t-1}\\ &=\sqrt{\overline{\alpha}_{t}}\bm{n}_{0}+\sqrt{1-\overline{\alpha}_{t}}\overline{\bm{\epsilon}}_{t},\end{split}$ (6)
where $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{t}\in(0,1)$ is the hyperparameter for adjusting the mean and variance of the $t$-th noise; $\epsilon_{t-1}$ and $\overline{\epsilon}_{t}$ are added noise all following the standard normal distribution, i.e., $\bm{\epsilon}_{t-1}$, $\overline{\bm{\epsilon}}_{t}\sim\mathcal{N}(0,\bm{I})$, where $\bm{I}$ denotes the identity matrix. When setting hyperparameters with $\overline{\alpha}_{T}\approx 0$ and $1-\overline{\alpha}_{T}\approx 1$, the perturbed data $\bm{n}_{T}$ with $T$ times Gaussian noises added will approximately follow the standard normal distribution, i.e., $\bm{n}_{T}\sim\mathcal{N}(0,\bm{I})$.
As the opposite of the forward diffusion process, the denoising process aims to recover the pure data from the noise distribution. Since forward diffusion and denoising are inverse Markov chains in the same representative form, the denoising transitions can be expressed as [12]
$q(\bm{n}_{t-1}|\bm{n}_{t})=\mathcal{N}(\bm{n}_{t-1};\bm{\mu}_{t},\bm{\sigma}_{t}^{2}\bm{I}).$ (7)
Generally, the denoising transitions in Eq. (7) are challenging to derive and compute directly. Thus, a noise neural network is trained to learn the transition distribution through the variational inference principle [13]. By the Bayes’ theorem, the learned transitions can be represented as
$q_{\theta}(\bm{n}_{t-1}|\bm{n}_{t},\bm{n}_{0})$ $=\mathcal{N}(\bm{n}_{t-1};\bm{\mu}_{\theta}(\bm{n}_{t},t),\bm{\sigma}_{t}^{2}\bm{I}),$ (8)
$\bm{\mu}_{\theta}(\bm{n}_{t},t)$ $=\frac{1}{\sqrt{\alpha_{t}}}(\bm{n}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{n}_{t},t)),$ (9)
$\bm{\sigma}_{t}^{2}$ $=\frac{(1-\alpha_{t})(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}},$ (10)
where $\bm{\epsilon}_{\theta}(\bm{n}_{t},t)$ is the noise neural network. The DM’s training aims to maximize the evidence lower bound (ELBO) on the log-likelihood of the original distribution $\log p(\bm{n})$ [14]. Given a trained noise network, new samples can be generated via sampling from estimated transitions, which can be expressed as
$\bm{n}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{n}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{n}_{t},t))+\bm{\sigma}_{t}\bm{\epsilon}_{t},$ (11)
where $\bm{\epsilon}_{t}\in\mathcal{N}(0,\bm{I})$, $t=T,T-1,\dots,1$, and $\bm{n}_{T}\in\mathcal{N}(0,\bm{I})$ is the initial noise. The sampling result $\bm{n}_{0}$ can be considered as the data from the original distribution $p(\bm{n})$, i.e., $\bm{n}_{0}\sim p(\bm{n})$.
III Proposed Signal Detection Method
In this section, we propose our Maximum Likelihood Estimation Diffusion Detector (MLEDD) framework for signal detection in near-field MIMO systems with unknown signal noise characteristics.
Figure 2: Architecture of the training and detection framework of MLEDD.
III-A Training of MLEDD
First, we collect the transmitted signal ${\bm{x}}$, received signal ${\bm{y}}$, and channel information ${\bm{A}}$ in the unknown noise scenario to establish a signal dataset $\mathcal{D}=\{({\bm{x}}^{i},{\bm{y}}^{i},{\bm{A}}^{i})\}_{i=1}^{|\mathcal{D}|}$. Then, a DM network $\bm{\epsilon}_{\theta}({\bm{n}}_{t},t)$ is trained to estimate the unknown signal noise distribution $p_{n}$, as shown in Fig. 2. In the training process, given the input $({\bm{x}}^{i},{\bm{y}}^{i},{\bm{A}}^{i})$, we perturb the corresponding noise ${\bm{n}}^{i}({\bm{x}}^{i},{\bm{y}}^{i},{\bm{A}}^{i})={\bm{y}}^{i}-{\bm{A}}^{i}{\bm{x}}^{i}$ via Gaussian noise to obtain the transition distribution $p_{n}({\bm{n}}^{i}_{t}|{\bm{n}}^{i}_{t-1})=\mathcal{N}({\bm{n}}^{i}_{t};\sqrt{\alpha_{t-1}}{\bm{n}}^{i}_{t-1},(1-\alpha_{t})\bm{I})$, where ${\bm{n}}^{i}_{0}={\bm{n}}^{i}$. To learn the reverse transition $q_{\theta}({\bm{n}}^{i}_{t-1}|{\bm{n}}^{i}_{t},{\bm{n}}^{i}_{0})$ in Eq. (8), we aim to maximize the ELBO via minimizing the KL-Divergence between $p_{n}$ and $q_{\theta}$. The loss function can be expressed as follows [11]:
$\begin{split}&l(\theta):=\mathbb{E}_{({\bm{x}}^{i},{\bm{y}}^{i},{\bm{A}}^{i})\in\mathcal{D}}\mathbb{E}_{t\in[0,T],\bm{\epsilon}\in\mathcal{N}(0,\bm{I})}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[||\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\overline{\alpha}_{t}}{\bm{n}}^{i}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon},t)||^{2}].\end{split}$ (12)
III-B Detecting of MLEDD
In the detecting process, given a received signal ${\bm{y}}$ and the channel matrix ${\bm{A}}$, our goal is to find the suitable signal ${\bm{x}}_{s}\in\mathcal{X}$, which is the MLE $\widehat{\bm{x}}_{MLE}$ in Eq. (5). However, it is challenging to derive a formal expression for the unknown noise distribution $p_{n}$ due to the difficulty of solving the integral calculation of the neural network $\bm{\epsilon}_{\theta}$.
Inspired by Diffusion Classifier (DC) [15], MLEDD detects the optimal signal by comparing the KL-divergence with perturbed distribution instead of directly obtaining it by denoising. We consider each potential signal $\bm{x}_{i}$ in the signal set $\mathcal{X}$ as a category and utilize MLE as the metric for classification. Via Bayes’ theorem, the MLE over label $\{\bm{x}_{i}\}$ can be written as follow:
$p({\bm{y},\bm{A}}|\bm{x}_{i})=\frac{p(\bm{x}_{i}|{{\bm{y},\bm{A}}})p({\bm{y}},{\bm{A}})}{p(\bm{x}_{i})}.$ (13)
Considering $p({\bm{y}},{\bm{A}})$ is constant and $p(\bm{x}_{i})$ is identical for all $i$ when there is no prior knowledge, we have
$\begin{split}p({\bm{y}},{\bm{A}}|\bm{x}_{i})\propto p(\bm{x}_{i}|{\bm{y}},{\bm{A}})=p_{n}(\bm{y}-\bm{A}\bm{x}_{i})=p_{n}(\bm{n}^{i}).\end{split}$ (14)
Algorithm 1 MLEDD
1:Timestep $T$, coefficient $\{\alpha_{t}\}_{t=1}^{T}$, training set $\mathcal{D}$, received signal $\bm{y}$, channel matrix $\bm{A}$, signal set $\mathcal{X}$, sampling pairs set $\{({t_{s}},{\bm{\epsilon}_{s}})\}_{s=1}^{S}$
2:repeat
3:    Choose $({\bm{x}}^{i},{\bm{y}}^{i},{\bm{A}}^{i})\in\mathcal{D}$, tsimilar-$t\sim$Uniform ($(\{1,\ldots,T\})$, and $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$
4:    Take gradient descent step on Eq. (12)
5:until converged
6:for all $\bm{x}_{i}\in\mathcal{X}$ do:
7:    Compute $E(\bm{x}_{i})$ in Eq. (16)
8:${\bm{x}}_{d}\leftarrow\arg\min_{\bm{x}_{i}\in\mathcal{X}}E(\bm{x}_{i})$
9:return ${\bm{x}}_{d}$
Similar to the training process of MLEDD, we use the ELBO of $\log q_{\theta}(\bm{n}^{i})$ in place of $\log p_{n}(\bm{n}^{i})$, i.e.,
$\begin{split}p_{n}(\bm{n}^{i})\propto\exp\{-\mathbb{E}_{t\in[0,T],\bm{\epsilon}\in\mathcal{N}(0,\bm{I})}[||\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{n}^{i}_{t},t)||^{2}]\},\end{split}$ (15)
where $\bm{\epsilon}_{\theta}({\bm{n}}_{t},t)$ is the trained DM model, and $\bm{n}^{i}_{0}={\bm{n}^{i}}=\bm{y}-\bm{A}\bm{x}_{i}$. Based on this, we define an error function $E(\bm{x}_{i})$ with the set of sampling pairs $\mathcal{S}=\{({t_{s}},{\bm{\epsilon}_{s}})\}_{s=1}^{S}$, which can be expressed as follows:
$E(\bm{x}_{i})=\frac{1}{S}\sum_{s=1}^{S}||{\bm{\epsilon}_{s}}-\bm{\epsilon}_{\theta}(\sqrt{\overline{\alpha}_{t_{i}}}{\bm{n}}^{i}_{0}+\sqrt{1-\overline{\alpha}_{t_{s}}}\bm{\epsilon}_{s},{t_{s}})||^{2}.$ (16)
For each possible signal vector $\bm{x}_{i}$ in $\mathcal{X}$, the detected signal ${\bm{x}}_{d}$ is the signal vector with the least error $E$ in Eq. (16). The overall MLEDD algorithm is shown as Algorithm 1.
III-C Network Architecture
In MLEDD, wireless signals are converted to the real-valued representation to express the complex-valued input equivalently, i.e.,
$\overbrace{\begin{bmatrix}Re(\bm{y})\\ Im(\bm{y})\\ \end{bmatrix}}^{\overline{\bm{y}}}=\overbrace{\begin{bmatrix}Re(\bm{A})&-Im(\bm{A})\\ Im(\bm{A})&Re(\bm{A})\\ \end{bmatrix}}^{\bm{\overline{A}}}\overbrace{\begin{bmatrix}Re(\bm{x})\\ Im(\bm{x})\\ \end{bmatrix}}^{\overline{\bm{x}}}+\overbrace{\begin{bmatrix}Re(\bm{n})\\ Im(\bm{n})\\ \end{bmatrix}}^{\overline{\bm{n}}},$ (17)
where $Re(\cdot)$ and $Im(\cdot)$ represent the real and imaginary part functions, respectively. Based on this, the received signal (Eq. (1)) can be reformulated as follow:
$\overline{\bm{y}}=\bm{\overline{A}\overline{x}}+\bm{\overline{n}}.$ (18)
As depicted in Fig. 2, the DM network utilizes two 1D convolution layers with a kernel size $k$ of 3 and rectified linear unit (ReLU) activation to progressively increase the number of channels to $C_{max}$. The embedded time $t$ is split into a scaling part and a bias part and then is connected to the output of convolution layers. Afterward, three 1D convolution layers are utilized by gradually decreasing the convolutional channels into the initial channel $C_{init}$. The same linear schedule of $\alpha_{t}$ is chosen as in [14].
IV Simulation Results
In this section, we conduct simulations to showcase the effectiveness of the proposed MLEDD detection framework. All experiments were conducted on an A100 Linux server using Python and the PyTorch package. The main program is based on the DM framework in [14].
IV-A Environment Set
We consider the number of MTs and antennas of the BS as $(M,N)=(4,5)$ in the near-field MIMO scenario. The signals transmitted have a carrier frequency of $f=28$ GHz, corresponding to a wavelength of $\lambda=0.0107$ m. We set the array spacing as $d=\lambda/2$. In this case, the Fresnel region is between $112\lambda$ and $2048\lambda$ [10]. The modulation scheme is quadrature phase shift keying (QPSK) with $P=4$. We assume the receiver has perfect knowledge of the channel state information (CSI).
[…]
IV-B Baseline Approach
We compare the proposed MLEDD with the following classical and learning-based baselines. The classical Euclidean distance-based MLE detector following $\bm{x}_{MLE}=\arg\min_{\bm{x}\in\mathcal{X}}||\bm{y}-\bm{A}\bm{x}||^{2}$ is denoted as “MLE”, which is equivalent to minimum distance decision rule when noise is Gaussian. Additionally, two model-based linear detectors based on MMSE with variance as 1 and Zero-Forcing (ZF) are denoted as “MMSE” and “ZF”, respectively [4]. A learning-based signal detection framework DetNet from [6] with 30 layers is represented as “DetNet”. Additionally, we compare the proposed MLEDD with another generative model-based method labeled “MANFE” with the same setting in [8]. Specifically, each flow step consists of an activation normalization layer, an invertible $1\times 1$ convolution, and an alternating affine coupling layer. All training approaches are trained for 500 epochs with an early stopping mechanism in place [14].
(a) BER vs $\alpha$ with SNR = 5 dB
(b) BER vs $\alpha$ with SNR = 25 dB
(c) BER vs SNR with $\alpha$ = 1.1
(d) BER vs SNR with $\alpha$ = 1.9
Figure 3: BER performance comparison curves for $4$ methods in $4\times 5$ near-field MIMO system.
IV-C Performance Evaluation
To evaluate the BER performance of the proposed MLEDD method and three baselines, we compare the relationship between the noise parameter $\alpha$ and BER at a given SNR, as well as the relationship between BER and SNR at a given parameter $\alpha$, shown in Fig. 3.
First, we examine the BER performance of various methods under SNRs of 5 and 25. Fig. 3(a) illustrates that in low SNR scenarios, the proposed MLEDD method outperforms or has competitive performance with other methods in most impulsive noise environments. Specifically, when $\alpha=1.1$, where noise is highly impulsive and more extreme values are more likely to occur, the MLEDD can significantly reduce the BER of MLE, DetNEt, and MANFE to about 22.9%, 22.3%, and 16.3% respectively. Additionally, when the impulsive noise approaches a Gaussian distribution with $\alpha=2$, the performance improvement becomes less significant. When $\alpha>1.8$, although the performance is not as good as that of DetNet, a traditional model-based AI approach, it still exhibits slightly lower error rates than the existing generative learning method MANFE. On the other hand, when SNR is 25 dB in Fig. 3(b), MLEDD achieves the best performance on all impulsive noise levels. Even compared with MANFE, which has the closest performance, MLEDD reduces the average BER to about 31.1%.
Then, we investigate the BER performance under different noise distributions where $\alpha$ is $1.1$ and $1.9$, respectively. We can observe from Fig. 3(c) that MLEDD has a stronger learning ability for complex unknown noise distribution under all SNR conditions. Compared with model-based methods, including numerical method, MLE, and learning method, DetNet, MLEDD, the BER of MLEDD is significantly reduced, and as the SNR increases, the reduction is more obvious. In particular, when SNR is 30 dB, MLEDD can reduce the BER by 88.3% and 93.3%, respectively. As for the generative model MANFE, MLEDD still achieves around 22.8% average performance improvement, indicating MLEDD can handle it well under any communication quality for complex noise distributions. In Fig. 3(d), it is demonstrated that even for noise close to a Gaussian distribution, MLEDD performs similarly to other methods, and MLEDD shows significant improvement as the SNR increases. Specifically, at 30 dB SNR, MLEDD outperforms MLE by 27.4% and MANFE by 11.7%.
Additionally, the online complexity for likelihood estimation methods MLE, MANFE, and MLEDD are $\mathcal{O}(MNP^{N})$, $\mathcal{O}((KM+MN)P^{N})$, and $\mathcal{O}((|\mathcal{S}|C_{max}kM+MN)P^{N})$, where $K$ represents the depth of the flow network and $k$ is the kernel size of MLEDD. As the number of MTs $M$ and the number of antennas $N$ increase, the complexity of the three methods is of the same magnitude. Although model-based detectors, including MMSE, ZF, and DetNet, have lower complexity, they require assumptions about the noise distribution, leading to worse performance in the case of unknown noise.
In summary, the proposed MLEDD algorithm has a strong ability to handle unknown distribution noise and has better performance than existing methods with the same magnitude of complexity as the noise becomes more complex. Moreover, in all near-field communication qualities, especially in scenarios with good communication quality, it still has comparable performance to other methods.
V Conclusion
In this letter, we present a near-field signal detection algorithm based on the generative diffusion model within unknown noise scenarios. The proposed detection method, MLEDD, utilized forward diffusion and the corresponding denoising process to learn the data feature and distribution characteristics of unknown noise. Based on maximum likelihood estimation, an error function via Bayes’ theorem was defined to detect the source signal. The simulation results demonstrated that the proposed MLEDD algorithm outperformed the classical numerical method and existing learning-based methods, especially when the unknown distribution is more intricate.